a. What is the probability that Z is between −1.51 and 1.81?

Given a standardized normal distribution (with a mean of 0 and a standard deviation of 1), complete parts (a) through (d) below. 


a. What is the probability that Z is between −1.51 and 1.81? 

The probability that Z is between −1.51 and 1.81 is _______.
Answer #1

Sol Given that, Z WN(0, 1) Here, we need to find P(-1.51EC1-81) p(-1.51ZZ<1.81). PlZc1.91) - PlZ2-1-51) from standard (from =
Answer #2

SOLUTION :


P(z between - 1.51 and 1.81)

= P(z ≤ 1.81) - P(z ≤ - 1.51)


From the cumulative ND table, we find the above cumulative probabilities :


= 0.9649 - 0.0655

= 0.8994 (ANSWER).


answered by: Tulsiram Garg